It takes a hamster 36 s to run 7 complete laps around his circular hamster track. A) What is the period of his run? s B) What is the frequency?
Answer #1

Solution) Number of laps , n = 7

Frequency , f = (Number of cycles)/s

f = 7/36 = 0.194 s^(-1)

Period , T = 1/f

T = 1/(0.194)

T = 5.15 s

(A) Period , T = 5.15 s

(B) Frequency , f = 0.194 s^(-1)